### v1

# EAS-SNN: End-to-End Adaptive Sampling and Representation for Event-based Detection with Recurrent Spiking Neural Networks

Abstract
Event cameras, with their high dynamic range and temporal resolution, are ideally suited for object detection, especially under scenarios with motion blur and challenging lighting conditions. However, while most existing approaches prioritize optimizing spatiotemporal representations with advanced detection backbones and early aggregation functions, the crucial issue of adaptive event sampling remains largely unaddressed. Spiking Neural Networks (SNNs), which operate on an event-driven paradigm through sparse spike communication, emerge as a natural fit for addressing this challenge. In this study, we discover that the neural dynamics of spiking neurons align closely with the behavior of an ideal temporal event sampler. Motivated by this insight, we propose a novel adaptive sampling module that leverages recurrent convolutional SNNs enhanced with temporal memory, facilitating a fully end-to-end learnable framework for event-based detection. Additionally, we introduce Residual Potential Dropout (RPD) and Spike-Aware Training (SAT) to regulate potential distribution and address performance degradation encountered in spike-based sampling modules. Through rigorous testing on neuromorphic datasets for event-based detection, our approach demonstrably surpasses existing state-of-the-art spike-based methods, achieving superior performance with significantly fewer parameters and time steps. For instance, our method achieves a 4.4% mAP improvement on the Gen1 dataset, while requiring 38% fewer parameters and three time steps. Moreover, the applicability and effectiveness of our adaptive sampling methodology extend beyond SNNs, as demonstrated through further validation on conventional non-spiking detection models.
Keywords:
Event-based Vision Spiking Neural Network Neuromorphic Computing1 Introduction
Event cameras, inspired by biological vision systems, present a transformative approach to dynamic visual sensing. Diverging from the conventional imaging model, where cameras capture scene illumination synchronously via a global shutter, event cameras operate on an asynchronous basis, logging logarithmic intensity changes at the pixel level. Each significant change in this context is referred to as an "event". This distinct operational paradigm grants event cameras extraordinary capabilities: exceptional temporal resolution (>10k fps), high dynamic range (>120 dB), and remarkably low power consumption (<10 mW) [37, 50, 15, 16]. These attributes render event cameras exceptionally suited for object detection under challenging environments characterized by rapid motion, extreme lighting conditions, or stringent energy constraints.
However, the inherent asynchronous and sparse characteristics of data captured by event cameras present significant obstacles for traditional dense neural networks, which are primarily designed to process synchronous, uniformly structured data. To reconcile the asynchronous event stream with the dense tensor formats utilized in ANNs, most existing frameworks incorporate a sampling-aggregation mechanism, also referred to as event representation or embedding [30, 18, 3, 19, 38]. This integration effectively alleviates the computational demands engendered by the high temporal resolution of event data. Building upon this foundation, researchers have delved into exploiting the unique spatiotemporal attributes of event data through refining event representations [31, 52, 32, 34, 8] and enhancing the design of detection architectures [41, 46, 32, 53, 49, 19, 22], thereby boosting network efficacy. Additionally, the aggregation module is integrated with downstream feature extraction components into a unified end-to-end training framework [18]. Regarding event-based sampling, several studies have endeavored to enhance the sampling module from a rule-based or search-based perspective [32, 38, 69, 5]. Nonetheless, these initiatives frequently suffer from the lack of high-level semantic information or are encumbered by artificial constraints and suboptimal search efficiency. So far, the quest for a differentiable, adaptive sampling module that facilitates end-to-end optimization within event-based detection frameworks remains unfulfilled.
On the other hand, SNNs, composed of biologically plausible spiking neurons, support asynchronous event-driven computation on neuromorphic architectures[1, 9, 45] and so present a high potential for fast inference and low power consumption. This positions SNNs as inherently compatible with the sparse nature of event data. Moreover, a spectrum of low-level representations [52, 31, 18] potentially allows for their encapsulation within a unified SNN layer [28]. Nevertheless, the challenge of training deep SNNs under long time steps persists, in spite of rapid advancements in spike-based learning algorithm [58, 51, 39, 62, 14, 35, 68, 55]. This makes it difficult to align the discretization window of SNNs with event response time, culminating in a discernible information temporal information loss. Therefore, the current event-based detection models based on SNNs exhibit a notable performance discrepancy when compared to their ANN counterparts.
To address these challenges, this work introduces ARSNN, a learnable adaptive sampling methodology with recurrent convolutional SNNs (Fig. 1a), enabling a comprehensive end-to-end learning paradigm for event representation and feature extraction. This methodology utilizes pixel-wise spiking neurons in the sampling module to capture the dynamic information delivered by events across receptive fields. Upon the information integration surpassing a predefined threshold, the events between the last and the current spike firing times are sampled as shown in Fig. 1b. By optimizing the spike firing time and membrane potential distribution, the sampling module can be trained and coordinated with downstream spike-based detectors, thereby greatly mitigating the temporal alignment loss between the raw event stream and SNNs. Moreover, the proposed SNN-based approach is optimally positioned for deployment on neuromorphic hardware, potentially curtailing overhead from data migration. The main contributions are outlined as follows:
1.
We first propose a learnable adaptive sampling method and realize an end-to-end event-based detection framework via SNNs that seamlessly integrates the processes of sampling, aggregation, and feature extraction.
2.
We develop the techniques of Residual Potential Dropout (RPD) and Spike-Aware Training (SAT) to mitigate the performance degradation problem in SNN-based sampling modules with the recurrent convolutional architecture.
3.
Empirical evaluations on neuromorphic datasets (N-Caltech 101 and Gen1) show the proposed method achieves superior performance over the state-of-the-art spike-based methods with less energy consumption compared to ANNs. Furthermore, the effectiveness of our adaptive sampling method is verified across dense detection models.
2 Related Work
2.1 Spiking Neural Networks
Spiking Neural Networks (SNNs) have raised substantial interest due to their asynchronous computation paradigm and promising energy efficiency. However, optimizing SNNs poses challenges, particularly due to the binary nature of spike firing. The advancements in backpropagation techniques for SNNs, including time-based [43, 64, 42, 63, 67] and activation-based methods [58, 51, 62, 20, 7, 35, 55], estimate gradients for firing time and spike generation successfully. Additionally, various spiking neuron models [14, 21, 60, 36, 6, 24] have been proposed to enhance network heterogeneity and performance. To maintain gradient flow in deep SNNs, specialized network modules [65, 11, 14] and architectures [13, 53, 27, 59, 66] have been developed considering temporal dynamics in spiking neurons. These collective efforts are narrowing the performance gap between SNNs and ANNs in various tasks, including object detection [61, 53], point cloud recognition [48], and automatic speech recognition [57].
2.2 Event-based Detection
In event-based detection, methodologies are bifurcated into two principal categories: dense and sparse representations. Dense representations typically convert event data into frame-based formats, subsequently processed by conventional dense detection algorithms [30, 3, 4, 19, 38]. This approach, while facilitating the utilization of pre-trained models within an event-based context [40], tends to significantly compress the inherent spatiotemporal richness of events, necessitating sophisticated event-to-frame conversion techniques. In pursuit of more effectively extracting the spatiotemporal information from event streams, researchers have integrated the recurrent connections [46, 32, 19] and attention mechanisms [54, 33]. Recent advancements [19, 38, 22] strive to minimize detection latency without substantial performance degradation, through the strategic redesign of network backbones and event aggregation modules, thus enhancing adaptability to scenarios demanding high-speed processing.
Conversely, sparse representations, embodied by Graph Neural Networks (GNNs) and SNNs, offer substantial energy efficiency due to their inherent sparse computational models. GNN models [34, 49] construct dynamic spatio-temporal graphs from event stream, updating node and edge states with incoming events. In parallel, SNN methodologies [8, 61, 53], characterized by the internal temporal dynamics and spike-driven computation, naturally complement the asynchronous nature of event stream. However, given the high temporal resolution of event cameras, event-based sampling is imperative to mitigate computational demands and enhance network generalization capabilities, even for sparse representation.
For event-based sampling, a spectrum of methodologies has emerged to refine rule-based strategies, taking into account the instantaneous rate of event generation [32, 29, 38]. Nevertheless, these strategies routinely encounter challenges associated with hyperparameter sensitivity, as well as the difficulty of maintaining an optimal balance between sparse and ambiguous textures in complex scenarios. Nikola et al. [69] have explored the optimization of sampling windows based on the Gromov-Wasserstein discrepancy from raw event data to their representations. Nonetheless, the technique has yet not fully integrated high-level semantic guidance. Cao et al. [5] have proposed the use of SNNs as a mechanism for global event triggering, advocating for a feedback-update mechanism that employs pseudo labels generated by downstream detectors to guide SNN training. Despite these advancements, the integration of ANNs and SNNs remains disjointed, bridged solely by pseudo labels, which introduces artificial constraints on the optimization landscape and compromises search efficiency. In this work, we investigate an adaptive event sampling strategy leveraging learnable SNNs, marking a step towards achieving comprehensive end-to-end event-based representation learning.
3 Methodology
3.1 Event-based Detection
Consider an event stream denoted as , where each event is characterized by its spatial coordinates and , a polarity value indicating the change in pixel intensity, and a timestamp marking the occurrence time of events (typically measured in microseconds). In event-based detection, annotations at a specific target time include object categories and associated ground truth bounding boxes , encapsulated as . Ideally, a real-time detector is capable of delivering predictions at the timestamp level:
Nonetheless, given the high temporal resolution of event cameras, it is challenging for event-based object detectors to process and interpret times per second. Moreover, the necessity for detecting moving objects at microsecond intervals is often not critical, as object motion typically occurs at a slower pace than the pixel response rate of event cameras [46]. Consequently, the current body of literature [32, 19, 53, 61] frequently incorporates a sampling operation and an aggregation operation to reduce the volume of sparse and asynchronous event data, thereby facilitating more efficient processing prior to the detection phase:
By employing fixed window sampling as an example, can be expressed as:
In this formula, denotes the global time window for sampling, and represents the interval of sampling slices. We designate the collection of events that meet the criteria for fixed window sampling within the -th step as . For the -th step sampling, a typical aggregation function, , based on event count histograms, is defined as follows:
Here, represents the Dirac delta function, which equals 1 only when . By employing this approach, the asynchronous and sparse event data are effectively down-sampled into a compact form that significantly alleviates the computational burden on the detector. However, existing sampling techniques often operate independently of downstream object detectors or multi-scale feature extractors, necessitating extensive hyperparameter tuning and showing limited adaptability.
3.2 Reccurent Convolutional SNNs for Adaptive Event Sampling
In this study, our goal is to develop a learnable sampling module, denoted as , utilizing SNNs to address the challenge of non-differentiable, end-to-end learning that constrains rule-based sampling methods.
The leaky-integrate-and-fire (LIF) model, popular in neuromorphic computing due to its minimal computational requirements, serves as the foundation for our approach. This model is mathematically expressed as follows:
[…]
This leaky-integrate-and-fire dynamic closely aligns with the desired properties of an event sampler. A subset of events should be selectively sampled as the accumulated information from the asynchronous event stream reaches a certain threshold. Consequently, we associate the spike firing in SNNs with event-based sampling, integrating this functionality into an end-to-end optimization framework as shown in Fig. 2.
Firstly, restricted by the poor support for sparse spike-driven computation on GPUs, we perform early aggregation using event count as shown in Eq. 4 with a small time slice and obtain the actual -step inputs for . It is worth noting that in general, the closer the timestamps of events are to the detection target timestamp , the greater the impact on the detection result at . Therefore, in this case, the input closer to in time is preferentially sent to . This operation is also consistent with the temporal information dynamics in the SNNs which has been observed that information becomes highly concentrated in earlier few timesteps [26]. Following early aggregation, a distinct spiking neuron is allocated to each spatial position and polarity , with each neuron exhibiting unique behavior by accumulating its spatially localized input over time.
where denotes a constant decay factor. Our empirical findings suggest that an effective sampler must possess an advanced temporal credit assignment capability to extract informative and discriminative substreams from the inherently noisy and fluctuating spatiotemporal event data. Consequently, we incorporate recurrent synapse connections, as depicted in the bottom left portion of Fig. 2, to enhance the temporal representational capacity of SNNs:
Here, the spike firing at the last time step contributes to the input current and decay factor at the current time step directly through recurrent synapse and . denotes the sigmoid function which normalizes the decay factor . We refer to SNNs after this modification as “RSNN” compared to Eq. 9, which effectively helps improve detection performance. When the spike is emitted with over-threshold membrane potential, the event slice between the last spike firing time and the current spike firing time is sampled, so as to achieve local sampling:
This approach infuses the adaptability into event sampling, driven by the inherent variability in firing times across spiking neurons located at distinct spatial coordinates. Notably, we adopt the hard reset to clear the membrane potential from the last sample. Following the event-based sampling phase, the subsequent challenge is to effectively integrate each sampled set into a coherent embedding. Empirically, we find that decoupling the sampling and the aggregation modules might hinder the training phase due to unstable gradient flow. Hence, in this work, we unify the sampling and the aggregation within a singular SNN. We directly employ the accumulation of membrane potential within the specified sampling window as the aggregation input for the downstream network:
Such a model incorporating the spike-driven sampling mechanism and the synaptic recurrent connection is referred to as the ‘ARSNN’ for the latter discussion.
3.3 Mitigating Degradation of Spike-based Sampling
In light of the adaptive sampling, it can be seamlessly integrated with the downstream detector for end-to-end training via backpropagation through time (BPTT) [56] with surrogate gradient functions [58, 51, 14]. Nonetheless, this aggregation of Eq. 12 implicitly optimizes the sampling module by relying solely on the aggregation of membrane potentials, which lacks information from the firing threshold for optimizing the spike firing timing. To elucidate, we leverage the gradients from the activation-based technique in SNN training to compute the update of firing time (detailed derivation provided in the Supplementary Material):
This reveals that the adjustment in firing time is solely dependent on the gradient of the membrane potential and the instantaneous change in potential at the firing moment. It disregards the relationship between the firing time update with the firing threshold . To counteract this issue, we introduce Spike-aware Training (SAT), which incorporates the firing spike into the aggregation function:
Given , this modification does not alter the forward computation of the aggregation function. However, it effectively rewires the gradient flow:
where is the surrogate gradient function used for circumventing the non-differential problem of the Heaviside function of Eq. 7. Here, represents the adjustment from threshold-based gradient redirection, enhancing the update of firing time especially when near the threshold , thereby achieving a more precise sampling away from critical points.
Moreover, the implementation of a spike-based sampling module introduces the dilemma of handling data after the final spike emission. An intuitive idea to address this involves incorporating the remaining information into the embedding, aiming to reduce information loss. However, this approach leads to a scenario where neurons, even without firing, accumulate substantial information for the downstream process. This inadvertently aggravates sampling degradation by diminishing the incentive for the sampling module to accurately learn spike timing. Hence, we propose the strategy of Residual Potential Dropout (RPD) to exclude all events following the last spike, enhancing the learning of precise firing moments. To illustrate the efficacy of RPD, we show the potential aggregation distribution with and without RPD. The results, as depicted in Fig. 3, generally reveal a more compact feature distribution with RPD. Moreover, employing SAT for gradient guidance (shown in the case of ARSNN+SAT), we observe that sampler training devoid of RPD tends to exhibit mode collapse towards the non-firing spectrum with concentrated negative membrane potentials.
4 Experiments
The effectiveness of the proposed approach is initially substantiated in Sec. 4.1. Subsequently, in Sec. 4.2, we compare our comprehensive framework with state-of-the-art event-based methods across benchmark datasets. Additionally, Sec. 4.3 explores the adaptability, hyper-parameter sensitivity, and sampling diversity, while Sec. 4.4 delves into its energy efficiency.
Implementation Details. Our experiments are executed on four RTX 3090 GPUs. We utilize SpikingJelly [12] to enhance the training efficiency of the spike-based backbone. The ADAM optimizer is uniformly applied across all tests, with learning rates of 1e-3 for Gen1 and 1e-4 for N-Caltech 101. In alignment with practices from [19, 22], data augmentation techniques including random Zoom-in and horizontal flipping are employed for both datasets. For GEN-1, event sampling is capped at 200ms prior to the timestamp of labeling to manage data processing duration. Unlike the hard reset used in the sampling module, downstream networks implement a soft reset strategy. Rather than employing a learning rate scheduler, we leverage model weight smoothing during training through the use of an exponential moving average with a momentum of .
Benchmark Datasets. The Neuromorphic-Caltech 101 (N-Caltech 101) dataset [44], captured by an ATIS event camera [47] traversing 101 object categories via three saccades, comprises 8,246 event streams, each embodying 300 ms of saccades alongside bounding box annotations. We partition this dataset into an 8:2 split for training and validation. The Prophesee Gen1 dataset [10] presents another rigorous neuromorphic benchmark, offering over 39 hours of automotive recordings with more than 228k car and 28k pedestrian annotations for event-based detection in diverse driving scenarios. Adhering to the original evaluation protocol, prediction boxes with either a side length below 20 pixels or a diagonal shorter than 30 pixels are excluded after non-maximum suppression.
4.1 Ablation Study
In this section, we conduct a comparative analysis to ascertain the efficacy of individual components within the proposed sampling module as detailed in Table 1. In addition to comparing with the common event count approach, we introduce two novel embedding strategies leveraging SNNs and RSNNs, predicated on potential aggregation as detailed in Eq. 12. The results show that the integration of recurrent synaptic connections significantly enhances network performance on both datasets, with observed improvements of 6.7% on N-Caltech 101 and 4.9% on Gen1. Notably, employing the SAT and RPD modules in isolation with adaptive sampling sometimes results in a performance decline compared to RSNN (58.6% vs. 63.4% in mAP on N-Caltech 101). Conversely, when SAT and RPD are synergistically combined within the SNN framework, there is a significant performance improvement, yielding increases of 3.0% and 1.2% in mAP over RSNN on the N-Caltech 101 and Gen1 datasets, respectively. This demonstrates the critical importance of integrating these components to achieve optimal network performance.
| Method | RPD | SAT | N-Caltech 101 | Gen1 | ||
| mAP | mAP | mAP | mAP | |||
| Event Count | - | - | 0.587 | 0.369 | 0.611 | 0.321 |
| SNN | - | - | 0.567 | 0.367 | 0.670 | 0.366 |
| RSNN | - | - | 0.634 | 0.412 | 0.719 | 0.399 |
| ARSNN | 0.600 | 0.398 | 0.687 | 0.373 | ||
| ARSNN | 0.586 | 0.391 | 0.714 | 0.393 | ||
| ARSNN | 0.592 | 0.376 | 0.714 | 0.398 | ||
| ARSNN | 0.664 | 0.437 | 0.731 | 0.409 | ||
4.2 Performance Comparison
Next, we evaluate our proposed method against existing techniques on the Gen1 and N-Caltech 101 datasets. In addition to the proposed ARSNN in the embedding module, we adopt PLIF [14] neurons with learnable decay and tdBN [65] across the other spiking layers to improve the overall efficacy of SNNs. The models, sized at 8.92M and 25.3M parameters, mirror the configurations of YOLOX-S and YOLOX-M [17] respectively, achieved through scaling the network width. As evidenced in Table 2, our technique outperforms the state-of-the-art spike-based model [53] in performance metrics (0.354 vs. 0.310 mAP) while operating with a reduced parameter (8.92 M vs. 14.40 M) under three time steps. Even compared with ANNs that feature well-designed attention mechanisms [19] and recurrent architectures [32, 46], our strategy maintains competitiveness and notably marks the first instance of achieving over 40% COCO mAP on Gen1 through sparse spike-driven computations, thereby enhancing energy efficiency as detailed in Sec. 4.4. With regards to N-Caltech 101, to the best of our knowledge, our approach represents the first spike-based solution to attain competitive performance, setting a new benchmark in the field.
In object detection, modules are designed for specific roles: backbones extract features, Feature Pyramid Networks (FPN) fuse multi-scale features, and detection heads handle coordinate regression and object recognition. We examined the effects of sparse spike representation by integrating a non-spiking detection head and FPN into a hybrid architecture. As shown in Table 2, a spiking FPN within a compact 8.92M network outperforms its non-spiking counterpart, indicating that spike-based architectures might offer regularization advantages for event-driven detection. However, spiking detection heads within constrained time steps appear to reduce network performance.
| Method | Embedding | Backbone | Head | Temporal | #Params | Timestep | ||
| RED [46] | Event Volum | CNN+RNN | SSD | Yes | 24.1M | - | - | 0.40 |
| RVT [19] | HIST. | Transformer+RNN | YOLOX | Yes | 18.5M | 21 | - | 0.475 |
| EGO-12 [69] | EGO-12 | Transformer | YOLOv6 | No | - | - | - | 0.504 |
| AEGNN [49] | Graph | GNN | YOLO | No | 20.0M | - | - | 0.163 |
| NVS-S [34] | Graph | GNN | YOLO | No | 0.9M | - | - | 0.086 |
| TAF [38] | CNN | CNN | YOLOX | No | 14.8M | - | - | 0.454 |
| ASTMNet [32] | 1D TCNN | CNN+RNN | SSD | Yes | >100M | 3 | - | 0.467 |
| MatrixLSTM [4] | LSTM | CNN | YOLOv3 | Yes | 61.5M | - | - | 0.310 |
| VC-DenseNet [8] | Voxel Cube | SNN | SSD | Yes | 8.2M | 5 | - | 0.189 |
| VC-MobileNet [8] | Voxel Cube | SNN | SSD | Yes | 12.64M | 5 | - | 0.174 |
| LT-SNN [25] | HIST. | SNN | YOLOv2 | Yes | - | - | - | 0.298 |
| KD-SNN [2] | HIST. | SNN | CenterNet | Yes | 12.97M | 5 | - | 0.229 |
| EMS-ResNet10 [53] | HIST. | SNN | YOLOv3 | Yes | 6.2M | 5 | 0.547 | 0.267 |
| EMS-ResNet18 [53] | HIST. | SNN | YOLOv3 | Yes | 9.34M | 5 | 0.565 | 0.286 |
| EMS-ResNet34 [53] | HIST. | SNN | YOLOv3 | Yes | 14.40M | 5 | 0.590 | 0.310 |
| TR-YOLO [61] | HIST. | SNN | YOLOv3 | Yes | 8.7M | 3 | 0.451 | - |
| EAS-SNN | ARSNN | SNN | YOLOX | Yes | 25.3M | 3 | 0.731 | 0.409 |
| YOLOX | Yes | 25.3M | 3 | 0.718 | 0.393 | |||
| YOLOX | Yes | 25.3M | 3 | 0.699 | 0.375 | |||
| EAS-SNN | ARSNN | SNN | YOLOX | Yes | 8.92M | 3 | 0.687 | 0.372 |
| YOLOX | Yes | 8.92M | 3 | 0.692 | 0.372 | |||
| YOLOX | Yes | 8.92M | 3 | 0.675 | 0.354 | |||
| denotes the model uses a spiking backbone, a spiking feature pyramid network, and a spiking detection head. | ||||||||
| denotes the model uses a spiking backbone, a spiking feature pyramid network, and a non-spiking detection head. | ||||||||
| denotes the model uses a spiking backbone, a non-spiking feature pyramid network, and a non-spiking detection head. | ||||||||
| Method | Embedding | Backbone | Head | Temporal | #Params | Timestep | mAP mAP | |
| YOLE [3] | Leaky Surface | CNN | YOLO | No | - | - | 0.398 | - |
| AEGNN [49] | Graph | GNN | YOLO | No | 20.0M | - | 0.595 | - |
| NVS-S [34] | Graph | GNN | YOLO | No | 0.9M | - | 0.346 | - |
| ASYNet [41] | HIST. | CNN | YOLO | No | - | - | 0.643 | - |
| ASYNet [41] | QUE. | CNN | YOLO | No | - | - | 0.615 | - |
| EAS-SNN | ARSNN | SNN | YOLO-X | Yes | 25.3M | 3 | 0.664 | 0.437 |
| EAS-SNN | ARSNN | SNN | YOLO-X | Yes | 8.92M | 3 | 0.538 | 0.338 |
4.3 Experimental Analysis
Scalability to Dense Neural Network:
In this section, our objective is to validate the efficacy of the proposed adaptive sampling technique within conventional dense neural networks. To this end, we configure to transmit only the initial aggregated potential to the subsequent networks. Fig. 4b displays the mAP training curves on the N-Caltech 101 dataset, comparing different methods. The combination of ARSNN with SAT and RPD demonstrates promising performance improvements, while ARSNN with SAT or RPD alone leads to declined performance, highlighting the synergistic benefits of our proposed modules. Additionally, introducing recurrent synaptic connections further enhances spike-based embedding performance. These results validate the adaptability of our sampling module for integration into dense neural networks.
Impact of Early Aggregation: We investigate the influence of early aggregation, as defined in Eq. 4, by modulating the aggregation step on Gen1 using spiking YOLOX-S. Here, the first three slices sampling from -step inputs are aggregated and fed sequentially into a spike-based backbone. The parameter indicates the temporal resolution of inputs, where a higher corresponds to a finer granularity. As depicted in Fig. 4a, optimal performance is generally observed when remains consistent between the training and testing phases. Notably, the model trained with , sustains over 30% mAP, despite a fourfold reduction in granularity. Overall, models trained with demonstrate considerable robustness, delivering promising performance under all tested conditions.
Distribution of Event Sampling: In this part, we employ the model trained under and examine the behavior of each neuron while using for inference. The sampling intervals of various spiking neurons are shown in the last column of Fig. 5, which demonstrates significant variability. The longest sample covers the full input, while the shortest sample only includes the input of one time step. Here, we disregard the neurons without spike firing, which typically make up due to the sparse event input. Additionally, compared to the event count representation in the first column, the image of the firing count in the second column offers a clearer estimation of the scene.
4.4 Energy Efficiency
While ANNs execute a fixed number of multiplication-and-accumulation (MAC) operations determined by their architecture, SNNs on neuromorphic hardware utilize input-conditioned spikes for sparse accumulation (AC) operations, enhancing energy efficiency. To assess the energy efficiency of SNNs, we compute the average AC operation count across Gen1, as detailed in Table 4. The energy cost for 32-bit floating-point addition (AC) and multiply-and-add (MAC) are quantified at and per operation, respectively, following [23]. The results show a significant reduction in energy consumption ( for YOLOX-M and for YOLOX-S) with spike-driven computation. It is worth noting that, for ANNs, we omit the energy consumption of the embedding phase, offering a comparatively conservative estimate of power savings. The detailed firing rates and more results across different models can be found in Supplementary Materials.
| Module | YOLOX-M | Spiking YOLOX-M | YOLOX-S | Spiking YOLOX-S | ||||||
| MAC(G) | Energy(mJ) | AC(G) | MAC(G) | Energy(mJ) | MAC(G) | Energy(mJ) | AC(G) | MAC(G) | Energy(mJ) | |
| Embedding | 0 | 0 | 0.02 | 1.63 | 7.52 | 0 | 0 | 0.02 | 1.63 | 7.52 |
| Backbone | 16.21 | 74.57 | 8.49 | 1.41 | 14.12 | 5.24 | 24.10 | 2.66 | 0.71 | 5.64 |
| FPN | 8.19 | 37.68 | 3.40 | 0 | 3.06 | 2.63 | 12.09 | 1.26 | 0 | 1.14 |
| Head | 11.33 | 52.10 | 3.78 | 0 | 3.40 | 5.04 | 23.17 | 1.79 | 0 | 1.61 |
| Total | 35.73 | 164.35(1) | 15.68 | 3.04 | 28.10(5.85) | 12.90 | 59.35(1) | 5.74 | 2.34 | 15.91() |
5 Conclusion
We introduce the Adaptive Sampling technique with Recurrent Spiking Neural Networks (ARSNN), enhanced by Residual Potential Dropout (RPD) and Spike-Aware Training (SAT) specifically to mitigate performance degradation in event-based sampling. Integrated with a spike-based architecture in the downstream neural network, our comprehensive framework, EAS-SNN, achieves superior performance relative to existing spike-based detection methods, delivering enhanced energy efficiency in comparison to ANNs. To the best of our knowledge, it is the first effort toward end-to-end optimization for event-based sampling and representation. Additionally, we demonstrate the versatility of our proposed sampling technique through its application to dense neural networks. Despite our exploration of the robustness across varying granularity of early aggregation, the ultimate goal is the development of a spike-based sampler that can precisely synchronize with the temporal granularity of raw event streams, measured in microseconds, without necessitating early aggregation. Our future work will focus on achieving this granularity, pending advancements in neuromorphic computing infrastructure.