### v2

# EAS-SNN: End-to-End Adaptive Sampling and Representation for Event-based Detection with Recurrent Spiking Neural Networks

Abstract
Event cameras, with their high dynamic range and temporal resolution, are ideally suited for object detection in scenarios with motion blur and challenging lighting conditions. However, while most existing approaches prioritize optimizing spatiotemporal representations with advanced detection backbones and early aggregation functions, the crucial issue of adaptive event sampling remains largely unaddressed. Spiking Neural Networks (SNNs), operating on an event-driven paradigm, align closely with the behavior of an ideal temporal event sampler. Motivated by this, we propose a novel adaptive sampling module that leverages recurrent convolutional SNNs enhanced with temporal memory, facilitating a fully end-to-end learnable framework for event-based detection. Additionally, we introduce Residual Potential Dropout (RPD) and Spike-Aware Training (SAT) to regulate potential distribution and address performance degradation encountered in spike-based sampling modules. Empirical evaluation on neuromorphic detection datasets demonstrates that our approach outperforms existing state-of-the-art spike-based methods with significantly fewer parameters and time steps. For instance, our method yields a 4.4% mAP improvement on the Gen1 dataset, while requiring 38% fewer parameters and only three time steps. Moreover, the applicability and effectiveness of our adaptive sampling methodology extend beyond SNNs, as demonstrated through further validation on conventional non-spiking models. Code is available at Github.
Keywords:
Event-based Vision Spiking Neural Network Neuromorphic Computing1 Introduction
Event cameras, inspired by biological vision systems, present a transformative approach to dynamic visual sensing. Diverging from the conventional imaging model, where cameras capture scene illumination synchronously via a global shutter, event cameras operate on an asynchronous basis, logging logarithmic intensity changes at the pixel level. Each significant change in this context is referred to as an "event". This distinct operational paradigm grants event cameras extraordinary capabilities: exceptional temporal resolution (>10k fps), high dynamic range (>120 dB), and remarkably low power consumption (<10 mW) [39, 54, 16, 17]. These attributes render event cameras exceptionally suited for object detection under challenging environments characterized by rapid motion, extreme lighting conditions, or stringent energy constraints.
However, the inherent asynchronous and sparse characteristics of data captured by event cameras present significant challenges for traditional dense neural networks, which are primarily designed to process synchronous, uniformly structured data. To bridge the asynchronous event stream with the dense tensor formats utilized in ANNs, most existing frameworks incorporate a sampling-aggregation mechanism, also referred to as event representation [32, 19, 3, 20, 40] or event embedding. This integration effectively alleviates the computational demands engendered by the high temporal resolution of event data. Building upon this foundation, researchers have delved into exploiting the unique spatiotemporal attributes of event data through refining event representations [33, 56, 34, 36, 8] and enhancing the design of detection architectures [43, 49, 34, 57, 53, 20, 24], thereby boosting network efficacy. Additionally, the aggregation module is integrated with downstream feature extraction components into a unified end-to-end training framework [19]. Regarding event-based sampling, several studies have endeavored to enhance the sampling module from a rule-based or search-based perspective [34, 40, 75, 5]. Nonetheless, these initiatives often suffer from the lack of high-level semantic information or are hindered by artificial constraints and suboptimal search efficiency. So far, the quest for a differentiable, adaptive sampling module that facilitates end-to-end optimization within event-based detection frameworks remains unfulfilled.
On the other hand, SNNs, composed of biologically plausible spiking neurons, support asynchronous event-driven computation on neuromorphic hardware[1, 9, 47] and so present a high potential for fast inference and low power consumption. This inherent compatibility with event data makes SNNs ideally suited for event-based vision tasks. Moreover, a spectrum of low-level representations [56, 33, 19] can be encapsulated within a unified SNN layer [30]. Nevertheless, the challenge of training deep SNNs under long time steps persists, despite rapid advancements in spike-based learning algorithm [63, 55, 67, 15, 37, 74, 59]. This makes it difficult to align the discretization window of SNNs with event response time, culminating in temporal information loss. Consequently, the current event-based detection models based on SNNs exhibit a notable performance gap when compared to their ANN counterparts.
To address these challenges, this work introduces ARSNN, a learnable adaptive sampling methodology with recurrent convolutional SNNs (Fig. 1a), enabling a comprehensive end-to-end learning paradigm for event representation and feature extraction. This methodology utilizes pixel-wise spiking neurons in the sampling module to capture the dynamic information delivered by events across receptive fields. Upon the information integration surpassing a predefined threshold, the events between the last and the current spike firing times are sampled as shown in Fig. 1b. By optimizing the spike firing time and membrane potential distribution, the sampling module can be trained and coordinated with downstream spike-based detectors, thereby greatly mitigating the temporal alignment loss between the raw event stream and SNNs. The main contributions are outlined as follows:
1.
We first propose a learnable adaptive sampling method and implement an end-to-end event-based detection framework via SNNs, integrating the processes of sampling, aggregation, and feature extraction.
2.
We develop the techniques of Residual Potential Dropout (RPD) and Spike-Aware Training (SAT) to mitigate performance degradation in SNN-based sampling modules with the recurrent convolutional architecture.
3.
Empirical evaluations on neuromorphic datasets (N-Caltech 101, Gen1, and 1Mpx) demonstrate that the proposed method outperforms state-of-the-art spike-based methods with less energy consumption compared to ANNs. Furthermore, the effectiveness of our adaptive sampling method is verified across dense detection models.
2 Related Work
2.1 Spiking Neural Networks
Spiking Neural Networks (SNNs) have raised substantial interest due to their asynchronous computation paradigm and promising energy efficiency. However, optimizing SNNs poses challenges due to the binary nature of spike firing. The advancements in backpropagation techniques for SNNs, including time-based [45, 69, 44, 68, 73] and activation-based methods [63, 55, 67, 22, 7, 37, 59], estimate gradients for firing time and spike generation successfully. Additionally, various spiking neuron models [15, 23, 65, 38, 6, 26] have been proposed to enhance network heterogeneity and performance. To maintain gradient flow in deep SNNs, specialized network modules [70, 12, 15] and architectures [14, 57, 29, 64, 71] have been developed considering temporal dynamics in spiking neurons. These collective efforts are narrowing the performance gap between SNNs and ANNs in various tasks, including object detection [60, 57, 66], action recognition [21, 41, 11], point cloud recognition [52], and automatic speech recognition [62].
2.2 Event-based Detection
In event-based detection, methodologies are bifurcated into two principal categories: dense and sparse representations. Dense representations typically convert event data into frame-based formats, subsequently processed by conventional detection algorithms [32, 3, 4, 20, 40]. This approach, while facilitating the use of pre-trained models within an event-based context [42], compress the inherent spatiotemporal richness of events, necessitating sophisticated event-to-frame conversion techniques. To extract the spatiotemporal information more effectively, researchers have integrated the recurrent connections [49, 34, 20] and attention mechanisms [58, 35]. Recent advancements [20, 40, 24] strive to minimize detection latency without substantial performance degradation, through the strategic redesign of network backbones and event aggregation modules.
Conversely, sparse representations, embodied by Graph Neural Networks (GNNs) and SNNs, offer substantial energy efficiency due to their inherent sparse computational models. GNN models [36, 53] construct dynamic spatio-temporal graphs from event stream, updating node and edge states with incoming events. In parallel, SNN methods [8, 66, 57], characterized by the internal temporal dynamics and spike-driven computation, naturally complement the asynchronous nature of event stream. However, given the high temporal resolution of event cameras, event-based sampling is imperative to mitigate computational demands and enhance network generalization capabilities, even for sparse representation.
For event-based sampling, some works have emerged to refine rule-based strategies by considering the instantaneous rate of event generation [34, 31, 40]. Nevertheless, these strategies frequently encounter challenges associated with hyperparameter sensitivity and the difficulty of maintaining an optimal balance between sparse and ambiguous textures in complex scenarios. Nikola et al. [75] explored the optimization of sampling windows based on the Gromov-Wasserstein discrepancy from raw event data to their representations. Nonetheless, this technique has yet not fully integrated high-level semantic guidance. Cao et al. [5] proposed using SNNs for global event triggering, advocating for a feedback-update mechanism that employs pseudo labels generated by downstream detectors to guide SNN training. Despite these advancements, the integration of ANNs and SNNs remains disjointed, bridged solely by pseudo labels, which introduces artificial constraints on the optimization landscape and compromises search efficiency. In this work, we investigate an adaptive event sampling strategy leveraging learnable SNNs, marking a step towards achieving comprehensive end-to-end event-based representation learning.
3 Methodology
3.1 Event-based Detection
Consider an event stream denoted as , where each event is characterized by its spatial coordinates and , a polarity value indicating the change in pixel intensity, and a timestamp marking the occurrence time of events (typically measured in microseconds). In event-based detection, annotations at a specific target time include object categories and associated ground truth bounding boxes , encapsulated as . Ideally, a real-time detector is capable of delivering predictions at the timestamp level:
Nonetheless, given the high temporal resolution of event cameras, it is challenging for event-based object detectors to process and interpret times per second. Moreover, the necessity for detecting moving objects at microsecond intervals is often not critical, as object motion typically occurs at a slower pace than the pixel response rate of event cameras [49]. Consequently, most works [34, 20, 57, 66] frequently incorporate a sampling operation and an aggregation operation to reduce the volume of sparse and asynchronous event data, thereby facilitating more efficient processing prior to the detection phase:
By employing fixed window sampling as an example, can be expressed as:
where denotes the global window for sampling, and represents the interval of sampling slices. We designate the index set that meets the criteria for fixed window sampling within the -th step as . For the -th step sampling, a typical aggregation function, , based on event count histograms, is defined as follows:
Here, represents the Dirac delta function, which equals 1 only when . By employing this approach, the asynchronous and sparse event data are effectively down-sampled into a compact form that significantly alleviates the computational burden on the detector. However, existing sampling techniques often operate independently of downstream object detectors , necessitating extensive hyperparameter tuning and showing limited adaptability.
3.2 Reccurent Convolutional SNNs for Adaptive Event Sampling
In this study, our goal is to develop a learnable sampling module (Fig. 2), denoted as , utilizing SNNs to address the challenge of non-differentiable, end-to-end learning that constrains rule-based sampling methods.
The Leaky-Integrate-and-Fire (LIF) model, popular in neuromorphic computing due to its minimal computational requirements. The discrete LIF equations, via the first-order Euler method with a time slice of , are given by:
Here, and denote the membrane potential and input current of the -th neuron in the -th layer, respectively, with as the membrane time constant. Upon receiving spike inputs, a neuron accumulates the input current into its membrane potential , which decays at a constant rate. When the membrane potential exceeds the threshold , a spike is emitted, resetting to . The LIF dynamic closely aligns with the desired properties of event-driven sampling, where events should be selectively sampled as the accumulated information from the asynchronous event stream reaches a threshold. Consequently, we propose to associate spike firing in SNNs with event-based sampling, integrating this functionality into an end-to-end optimization framework.
Firstly, restricted by the poor support for sparse spike-driven computation on GPUs, we perform early aggregation using event count as shown in Eq. 4 with a small time slice and obtain the actual -step inputs for . Following early aggregation, a distinct spiking neuron is allocated to each spatial position and polarity , with each neuron exhibiting unique behavior by accumulating its spatially localized input over time:
where denotes a constant decay factor. Our empirical findings suggest that an effective sampler must possess an advanced temporal credit assignment capability to extract informative and discriminative substreams from the inherently noisy and fluctuating spatiotemporal event data. Therefore, we incorporate recurrent synaptic connections, as depicted in the bottom left portion of Fig. 2, to enhance the temporal representational capacity of SNNs:
[…]
This approach optimizes event sampling by converting it into learning proper firing times across spiking neurons at distinct spatial coordinates. After event-based sampling, the remained challenge is to effectively integrate each sampled set into a coherent embedding. Empirically, we find that decoupling the sampling and the aggregation modules can hinder training due to unstable gradient flow. Thus, we unify the sampling and the aggregation within a single SNN, using the potential accumulation within the specified sampling window as the aggregation input for the downstream network:
Notably, we adopt the hard reset to clear the membrane potential from the previous sample. Such a model incorporating the spike-driven sampling mechanism and the synaptic recurrent connection is referred to as the ‘ARSNN’ for the latter discussion.
3.3 Mitigating Degradation of Spike-based Sampling
In light of the adaptive sampling, it can be integrated with the downstream detector for end-to-end training via backpropagation through time (BPTT) [61] with surrogate gradient functions [63, 55, 15]. However, this aggregation method in Eq. 9 implicitly optimizes the sampling module by relying solely on the aggregation of membrane potentials, which lacks information from the firing threshold for optimizing the spike firing timing. Formally, we leverage the gradients from the activation-based technique in SNN training to estimate the update of firing time (detailed derivation provided in the Supplementary Material):
This reveals that the adjustment in firing time depends solely on the gradient of the membrane potential and the instantaneous change in potential at the firing moment. It disregards the relationship between the firing time update and the firing threshold . To address this issue, we introduce Spike-aware Training (SAT), which incorporates the firing spike into the aggregation function as follows:
Given , this modification does not alter the forward computation of the aggregation function. However, it effectively rewires the gradient flow:
where is the surrogate gradient function used to circumvent the non-differential problem of the Heaviside function in Eq. 5. Here, represents the adjustment from threshold-based gradient redirection, enhancing the update of firing time especially when near the threshold , thereby achieving more precise sampling away from critical points. Moreover, the implementation of a spike-based sampling module introduces the dilemma of handling data after the final spike firing. An intuitive idea is to incorporate the remaining information into the embedding to reduce information loss. However, this leads to a scenario where neurons, even without firing, accumulate substantial information for representation. This inadvertently causes sampling degradation by diminishing the incentive for the sampling module to learn accurate spike timing. Hence, we propose the strategy of Residual Potential Dropout (RPD) to exclude all events following the last spike, enhancing the learning of precise firing moments. To illustrate the efficacy of RPD, we show the potential aggregation distribution with and without RPD. The results in Fig. 3 reveal a more compact feature distribution with RPD. Moreover, employing SAT for gradient guidance (as in the case of ARSNN+SAT), we observe that sampler training without RPD tends to exhibit mode collapse towards the non-firing spectrum with concentrated negative membrane potentials.
4 Experiments
The effectiveness of the proposed approach is first validated in Sec. 4.1. In Sec. 4.2, we compare our framework with state-of-the-art event-based methods across benchmark datasets. Sec. 4.3 explores adaptability, hyper-parameter sensitivity, and sampling diversity, while Sec. 4.4 examines energy efficiency.
Implementation Details. The ADAM optimizer is applied with learning rates of 1e-3 for Gen1, 1e-4 for N-Caltech 101, and 5e-4 for 1Mpx. Consistent with [20, 24], data augmentation techniques, including random Zoom-in and horizontal flipping, are employed for all datasets. For Gen-1, event sampling is capped at 200ms prior to the timestamp of labeling. For 1Mpx, we preprocess the event stream following [20], while using a subsequence of length 4 for BPTT training. Unlike the hard reset used in the sampling module, downstream networks implement a soft reset strategy. Instead of employing a learning rate scheduler, we leverage model weight smoothing during training through the use of an exponential moving average with a momentum of . Additionally, we utilize SpikingJelly [13] to enhance the training efficiency of the spike-based backbone.
Benchmark Datasets. The Neuromorphic-Caltech 101 (N-Caltech 101) dataset [46], captured by an ATIS event camera [50], comprises 8,709 event streams, each containing 300 ms of saccades with bounding box annotations and object contours. We partition this dataset into an 8:2 split for training and validation. The Prophesee Gen1 dataset [10] offers over 39 hours of automotive recordings, annotated with more than 228k car and 28k pedestrian labels, facilitating event-based detection in diverse driving scenarios. The 1Mpx dataset [49] features high-resolution driving recordings (1280×720) across various scenarios, including over 24 million labeled bounding boxes of cars, pedestrians, and two-wheelers, annotated at 60Hz. Following the original evaluation protocol, we exclude prediction boxes with a side length below 10 (20) pixels or a diagonal shorter than 30 (60) pixels after non-maximum suppression for the Gen1 (1Mpx) dataset.
4.1 Ablation Study
In this section, we conduct a comparative analysis to ascertain the efficacy of individual components within the proposed sampling module, as detailed in Table 2. In addition to comparing with Event Count used in early aggregation, we introduce two novel embedding strategies leveraging SNNs and RSNNs, predicated on potential aggregation. The results show that the integration of recurrent synaptic connections significantly enhances network performance on both datasets, with observed improvements of 6.7% on N-Caltech 101 and 4.9% on Gen1. Notably, employing the SAT and RPD modules in isolation with adaptive sampling sometimes results in a performance decline compared to RSNN (58.6% vs. 63.4% in mAP50 on N-Caltech 101). Conversely, when SAT and RPD are synergistically combined within the SNN framework, there is a significant performance improvement, yielding increases of 3.0% and 1.2% in mAP50 over RSNN on the N-Caltech 101 and Gen1 datasets, respectively. This demonstrates the critical importance of integrating these components to achieve optimal network performance. Furthermore, Table 2 compares our approach with established event representations such as Voxel Grid [72], Time Surface [56], and Voxel Cube [8]. It shows that the proposed adaptive method outperforms these alternatives on both datasets. Moreover, the framework can incorporate various event representations at the early aggregation stage for better performance.
Method | RPD | SAT | N-Caltech 101 | Gen1 | |||
mAP50 | mAP50:95 | mAP50 | mAP50:95 | ||||
Event Count | - | - | 0.587 | 0.369 | 0.611 | 0.321 | |
SNN | - | - | 0.567 | 0.367 | 0.670 | 0.366 | |
RSNN | - | - | 0.634 | 0.412 | 0.719 | 0.399 | |
ARSNN | 0.600 | 0.398 | 0.687 | 0.373 | |||
ARSNN | 0.586 | 0.391 | 0.714 | 0.393 | |||
ARSNN | 0.592 | 0.376 | 0.714 | 0.398 | |||
ARSNN | 0.664 | 0.437 | 0.731 | 0.409 |
Method | N-Caltech 101 | Gen1 | |||
mAP50 | mAP50:95 | mAP50 | mAP50:95 | ||
Event Count | 0.587 | 0.369 | 0.611 | 0.321 | |
Voxel Grid | 0.598 | 0.391 | 0.515 | 0.247 | |
Time Surface | 0.497 | 0.288 | 0.654 | 0.354 | |
Voxel Cube | 0.589 | 0.401 | 0.651 | 0.340 | |
SNN | 0.567 | 0.367 | 0.670 | 0.366 | |
RSNN | 0.634 | 0.412 | 0.719 | 0.399 | |
ARSNN | 0.664 | 0.437 | 0.731 | 0.409 |
Method | Representation | Backbone | Head | Temporal | #Params | Timestep | ||
RED [49] | Event Volum | CNN+RNN | SSD | Yes | 24.1M | - | - | 0.40 |
RVT [20] | HIST. | Transformer+RNN | YOLOX | Yes | 18.5M | 21 | - | 0.472 |
ERGO-12 [75] | ERGO-12 | Transformer | YOLOv6 | No | - | - | - | 0.504 |
AEGNN [53] | Graph | GNN | YOLO | No | 20.0M | - | - | 0.163 |
NVS-S [36] | Graph | GNN | YOLO | No | 0.9M | - | - | 0.086 |
TAF [40] | CNN | CNN | YOLOX | No | 14.8M | - | - | 0.454 |
ASTMNet [34] | 1D TCNN | CNN+RNN | SSD | Yes | >100M | 3 | - | 0.467 |
MatrixLSTM [4] | LSTM | CNN | YOLOv3 | Yes | 61.5M | - | - | 0.310 |
VC-DenseNet [8] | Voxel Cube | SNN | SSD | Yes | 8.2M | 5 | - | 0.189 |
VC-MobileNet [8] | Voxel Cube | SNN | SSD | Yes | 24.26M | 5 | - | 0.147 |
LT-SNN [27] | HIST. | SNN | YOLOv2 | Yes | - | - | - | 0.298 |
KD-SNN [2] | HIST. | SNN | CenterNet | Yes | 12.97M | 5 | - | 0.229 |
EMS-ResNet10 [57] | HIST. | SNN | YOLOv3 | Yes | 6.2M | 5 | 0.547 | 0.267 |
EMS-ResNet18 [57] | HIST. | SNN | YOLOv3 | Yes | 9.34M | 5 | 0.565 | 0.286 |
EMS-ResNet34 [57] | HIST. | SNN | YOLOv3 | Yes | 14.40M | 5 | 0.590 | 0.310 |
TR-YOLO [66] | HIST. | SNN | YOLOv3 | Yes | 8.7M | 3 | 0.451 | - |
EAS-SNN (M) | ARSNN | SNN | YOLOX⋄ | Yes | 25.3M | 3 | 0.731 | 0.409 |
YOLOX∗ | Yes | 25.3M | 3 | 0.718 | 0.393 | |||
YOLOX† | Yes | 25.3M | 3 | 0.699 | 0.375 | |||
EAS-SNN (S) | ARSNN | SNN | YOLOX⋄ | Yes | 8.92M | 3 | 0.687 | 0.372 |
YOLOX∗ | Yes | 8.92M | 3 | 0.692 | 0.372 | |||
YOLOX† | Yes | 8.92M | 3 | 0.675 | 0.354 | |||
denotes the model uses a spiking backbone, a spiking feature pyramid network, and a spiking detection head. | ||||||||
denotes the model uses a spiking backbone, a spiking feature pyramid network, and a non-spiking detection head. | ||||||||
denotes the model uses a spiking backbone, a non-spiking feature pyramid network, and a non-spiking detection head. |
Method | Representation | Backbone | Head | Temporal | #Params | Timestep | mAP50 mAP50:95 | |
YOLE [3] | Leaky Surface | CNN | YOLO | No | - | - | 0.398 | - |
AEGNN [53] | Graph | GNN | YOLO | No | 20.0M | - | 0.595 | - |
NVS-S [36] | Graph | GNN | YOLO | No | 0.9M | - | 0.346 | - |
ASYNet [43] | HIST. | CNN | YOLO | No | - | - | 0.643 | - |
ASYNet [43] | QUE. | CNN | YOLO | No | - | - | 0.615 | - |
EAS-SNN(M) | ARSNN | SNN | YOLO-X | Yes | 25.3M | 3 | 0.664 | 0.437 |
EAS-SNN(S) | ARSNN | SNN | YOLO-X | Yes | 8.92M | 3 | 0.538 | 0.338 |
4.2 Performance Comparison
We evaluate our proposed method against existing models on the Gen1, N-Caltech 101 and 1Mpx datasets. To improve the overall efficacy, we adopt PLIF [15] neurons with learnable decay and tdBN [70] in the downstream detector. The models, with sizes of 8.92M (SYOLOX-S) and 25.3M (SYOLOX-M) parameters, mirror the network architectures of YOLOX [18]. As shown in Table 3, our technique outperforms the state-of-the-art spike-based method [57] (0.354 vs. 0.310 mAP50:95) while using fewer parameters (8.92 M vs. 14.40 M) under three time steps. Even compared with ANNs that feature well-designed attention mechanisms [20] and recurrent architectures [34, 49], our method remains competitive and notably achieves over 40% mAP50:95 on Gen1 through sparse spike-driven operations, thereby enhancing energy efficiency as detailed in Sec. 4.4. Regarding N-Caltech 101 (Table 4), our approach improves the detection performance with spike-based representation. On the 1Mpx dataset, the largest event-based detection dataset, the proposed method achieves the first spike-based results, demonstrating remarkable performance using SNNs, as depicted in Table 5.
In object detection, each module has a specific role: backbones extract features, Feature Pyramid Networks (FPN) fuse multi-scale features, and detection heads handle coordinate regression and object recognition. We explored the impact of sparse spike representation by integrating a non-spiking detection head and FPN into a hybrid architecture. As shown in Table 3, a spiking FPN within a compact 8.92M network outperforms its non-spiking counterpart, indicating that spike-based architectures might offer regularization benefits for event-based detection. However, spiking detection heads within constrained time steps appear to reduce network performance.
Method | Representation | Backbone | Head | Temporal | #Params | Timestep | ||
RED [49] | Event Volum | CNN+RNN | SSD | Yes | 24.1M | - | - | 0.430 |
RVT [20] | HIST. | Transformer+RNN | YOLOX | Yes | 18.5M | 21 | - | 0.474 |
ERGO-12 [75] | ERGO-12 | Transformer | YOLOv6 | No | - | - | - | 0.406 |
ERGO-12 [75] | HIST. | Transformer | YOLOv6 | No | - | - | - | 0.327 |
E2VID [51] | Reconstruction | CNN | RetinaNet | No | - | - | - | 0.250 |
ASTMNet [34] | 1D TCNN | CNN+RNN | SSD | Yes | >100M | 3 | - | 0.483 |
Events+YOLOv3 [28] | HIST. | CNN | YOLOv3 | No | - | - | - | 0.346 |
EAS-SNN (M) | ARSNN | SNN | YOLOX | Yes | 25.3M | 3 | 0.651 | 0.362 |
4.3 Experimental Analysis
Scalability to Dense Neural Network:
In this section, our objective is to validate the efficacy of the proposed adaptive sampling technique within conventional dense neural networks. Here, we only transmit the initial aggregated potential to subsequent networks. Fig. 4b displays the mAP50 training curves on the N-Caltech 101 dataset, comparing different methods. The combination of ARSNN with SAT and RPD demonstrates promising performance improvements, while ARSNN with SAT or RPD alone leads to declined performance, highlighting the synergistic benefits of our proposed modules. Additionally, recurrent synaptic connections further improve network performance.
Impact of Early Aggregation: We investigate the influence of early aggregation, as defined in Eq. 4, by modulating the aggregation step on Gen1 using spiking YOLOX-S. The first three slices sampled from -step inputs are aggregated and fed sequentially into a spike-based backbone. The parameter indicates the temporal resolution of inputs, where higher corresponds to finer granularity. As depicted in Fig. 4a, optimal performance is generally observed when remains consistent between the training and testing phases. Notably, the model trained with , sustains over 30% mAP50:95, despite a fourfold increase in granularity. Overall, models trained with demonstrate considerable robustness, delivering promising performance under all tested conditions.
Distribution of Event Sampling: In this part, we employ the model trained under and examine the behavior of each neuron while using for inference. The sampling intervals of various spiking neurons are shown in the last column of Fig. 5, which demonstrates significant variability. The longest sample covers the full input, while the shortest sample only includes the input of one time step. Here, we disregard the neurons without spike firing, which typically make up due to the sparse event input. Additionally, compared to the event count representation in the first column, the image of the firing count in the second column offers a clearer estimation of the scene. The qualitative results of adaptive sampling are further presented in Fig. 6. The top row shows event count image without adaptive sampling, under blurred and sparse textures. The bottom row demonstrates the improvement in information retention and noise reduction with adaptive sampling in the same contexts.
4.4 Energy Efficiency and Inference Speed
While ANNs execute dense multiplication-and-accumulation (MAC) operations determined by their architecture, SNNs on neuromorphic hardware utilize input-conditioned spikes for sparse accumulation (AC) operations, thereby enhancing energy efficiency. To evaluate the energy efficiency of SNNs, we compute the average AC operations across Gen1, as detailed in Table 6. According to [25], we estimate the energy cost for 32-bit floating-point AC and MAC as and per operation, respectively. The results show a significant reduction in energy consumption with spike-driven computation, achieving a reduction for YOLOX-M and for YOLOX-S. It is worth noting that, for ANNs, we omit the energy consumption of the embedding phase, offering a comparatively conservative estimate of power savings. Detailed firing rates and additional results for different models are provided in Supplementary Materials.
Furthermore, we compared the frames per second (FPS) performance of our model with other state-of-the-art methods on the Gen1 dataset. Our model achieves competitive detection speeds of 54.35 FPS, compared to 59.88 FPS for RED [49], 28.09 FPS for ASTMNet [34], 106.38 FPS for RVT-T [20], and 59.52 FPS for GET [48]. This demonstrates the significant potential for higher efficiency on neuromorphic chips, which support sparse spike-driven operations.
Module | YOLOX-M | Spiking YOLOX-M | YOLOX-S | Spiking YOLOX-S | ||||||
MAC(G) | Energy(mJ) | AC(G) | MAC(G) | Energy(mJ) | MAC(G) | Energy(mJ) | AC(G) | MAC(G) | Energy(mJ) | |
Embedding | 0 | 0 | 0.02 | 1.63 | 7.52 | 0 | 0 | 0.02 | 1.63 | 7.52 |
Backbone | 16.21 | 74.57 | 8.49 | 1.41 | 14.12 | 5.24 | 24.10 | 2.66 | 0.71 | 5.64 |
FPN | 8.19 | 37.68 | 3.40 | 0 | 3.06 | 2.63 | 12.09 | 1.26 | 0 | 1.14 |
Head | 11.33 | 52.10 | 3.78 | 0 | 3.40 | 5.04 | 23.17 | 1.79 | 0 | 1.61 |
Total | 35.73 | 164.35(1) | 15.68 | 3.04 | 28.10(5.85) | 12.90 | 59.35(1) | 5.74 | 2.34 | 15.91() |
5 Conclusion
We introduce an adaptive sampling method utilizing recurrent spiking neural networks (ARSNN), enhanced by Residual Potential Dropout (RPD) and Spike-Aware Training (SAT) specifically to mitigate performance degradation in event-based sampling. Integrated with a spike-based architecture in the downstream neural network, our comprehensive framework, EAS-SNN, outperforms existing spike-based detection methods and offers superior energy efficiency compared to ANNs. To the best of our knowledge, it is the first effort toward end-to-end optimization for event-based sampling and representation. Additionally, we demonstrate the versatility of our proposed sampling technique through its application to dense neural networks. Despite exploring robustness across varying granularities of early aggregation, the ultimate goal is to develop a spike-based sampler that can precisely synchronize with the temporal granularity of raw event streams, measured in microseconds, without necessitating early aggregation. Our future work will focus on achieving this granularity, pending advancements in neuromorphic computing infrastructure.
Acknowledgements
This work was supported by National Natural Science Foundation of China under Grant 62236007, and the Key R&D Program of Zhejiang under Grant 2023C03001 and 2022C01048.